# LT3SD: Latent Trees for 3D Scene Diffusion

### v1

@@ C.2 Coarse-to-Fine Refinement
For large-scale 3D generation, generating patches autoregressively can be highly time-consuming. In the 3D scene depicted in Fig. 4, our full model initially generates the scene patch-by-patch, as outlined in Eq. 5, focusing on capturing the coarse structure. We then apply a coarse-to-fine refinement process, adding high-frequency details using the parallel algorithm, as described in Eq. 6. In the ablation study (Tab. 3), we replaced the parallel algorithm with the autoregressive one Eq. 5 to generate the finer levels. The results show that this approach uses 2.5×2.5$\times$2.5 × inference time, compared to our full model. This demonstrates that the batch-wise coarse-to-fine approach significantly reduces overall inference time for large-scale 3D scene generation, in contrast to the purely patch-by-patch method. Furthermore, our approach can be accelerated using multi-GPU setups, a capability that previous works [54, 32], which employ naive autoregressive patch-wise outpainting, do not support.

### v2

@@ B.1 Metrics
B.2 Baselines
To train XCube on the FRONT-3D dataset, we follow the open-sourced training scripts in two stages: VAEs and diffusion models. Each stage contains two levels: the coarse level compressing sparse scene patches of shape [128, 128, 64] to dense features of shape [32, 32, 16]; the fine level compressing sparse scene patches of shape [256, 256, 128] to sparse features of shape [128, 128, 64]. The resolution and voxel size of the finest level are determined by the limitations of the GPU memory (NVIDIA RTX A6000) and to remain consistent with other baselines. Since the first two stages only predict voxel occupancy, we utilize the pretrained refinement network of XCube-Objaverse to predict TUDF by making the voxel sizes and the origins of the grids the same.
B.3 Additional Comparisons
We provide additional qualitative comparison results in Fig. 10.
Appendix C Ablation Study
@@ C.2 Coarse-to-Fine Refinement
For large-scale 3D generation, generating patches autoregressively can be highly time-consuming. In the 3D scene depicted in Fig. 4, our full model initially generates the scene patch-by-patch, as outlined in Eq. 5, focusing on capturing the coarse structure. We then apply a coarse-to-fine refinement process, adding high-frequency details using the parallel algorithm, as described in Eq. 6. Fig. 4 shows that our highest resolution synthesis notably improves details in object structures such as chairs and pillows, compared to the coarser resolution shown on the left. In the ablation study (Tab. 3), we replaced the parallel algorithm with the autoregressive one Eq. 5 to generate the finer levels. The results show that this approach uses 2.5×2.5$\times$2.5 × inference time, compared to our full model. This demonstrates that the batch-wise coarse-to-fine approach significantly reduces overall inference time for large-scale 3D scene generation, in contrast to the purely patch-by-patch method. Furthermore, our approach can be accelerated using multi-GPU setups, a capability that previous works [57, 35], which employ naive autoregressive patch-wise outpainting, do not support.